OpenStudy (anonymous):

Any one in Algebra 1 Foundations A? Could use some help!

OpenStudy (shamil98):

Instead, just ask your question(s) regarding the topic.

OpenStudy (anonymous):

Their are over twenty questions I cant understand!

OpenStudy (shamil98):

Post them separately, one by one. I can try to help with your first one if you post it.

OpenStudy (anonymous):

alright thanks one sec

OpenStudy (anonymous):

What is the simplified form of the expression \[\sqrt{1/144}\]

OpenStudy (shamil98):

Do you know what a square root is? It is basically asking "What two numbers multiplied gives this number?"

OpenStudy (shamil98):

We can also write that as. \[\huge \frac{ \sqrt{1} }{ \sqrt{144} }\] this make it simpler?

OpenStudy (anonymous):

so it would be 1/12 right

OpenStudy (shamil98):

Yep

OpenStudy (anonymous):

Cool thanks!